Solve:
(1+x)² d²y/dx² + (1+x) dy/dx + y = 4 cos ln(1+x)
1
Expert's answer
2019-08-23T10:11:29-0400

"dy\/dx=y'"

"d^2y\/dx^2=y''"

"(1+x)² y'' + (1+x) y' + y = 4 \cos (\ln(1+x))"

Let's introduce the change of variable:

"1+x=e^t"

"t=\ln(1+x)"

"dt\/dx=1\/(1+x)=e^{-t}"

"\dot {y}=dy/dt"

"\ddot {y}=d^2y/dt^2"

"y'=\dot{y} e^{-t}"

"y''=(\ddot{y} -\dot{y})e^{-2t}"

So,

"\ddot{y}+\dot{y}+y=4\cos t"

We obtained nonhomogeneous linear differential equation of the second order. The general solution of this nonhomogeneous equation is the sum of the general solution "y_0" of the related homogeneous equation and a particular solution "y_1" of the nonhomogeneous equation: "y=y_0+y_1".

Related homogeneous differential equation:

"\ddot{y}+\dot{y}+y=0"

Its characteristic polynomial is "r^2+r+1". Discriminant "D=1^2-4 \cdot 1 \cdot 1=-3". So, the characteristic polynomial has two complex conjugate roots "\frac{-1±i \sqrt{3}}{2 \cdot 1}=-1/2±i \sqrt{3}/2" and the general solution of homogeneous differential equation is "y_0=e^{-t/2}(C_1 \cos (\sqrt{3}/2 \cdot t) + C_2 \sin (\sqrt{3}/2 \cdot t))".

Particular solution of the nonhomogeneous equation can be written in the form "y_1=A \cos t+B \sin t", where A and B - undetermined coefficients.

"\dot{y_1}=-A \sin t+B \cos t"

"\ddot{y_1}=-A \cos t - B \sin t"

Let's substitute these values into "\ddot{y}+\dot{y}+y=4\cos t":

"-A \cos t - B \sin t - A \sin t+B \cos t + A \cos t + B \sin t = 4 \cos t"

"B \cos t - A \sin t = 4 \cos t"

"\begin{cases}
A=0 \\
B=4
\end{cases}"

"y_1=4 \sin t"

So, "y=y_0+y_1=e^{-t/2}(C_1 \cos (\sqrt{3}/2 \cdot t)+ C_2 \sin (\sqrt{3}/2 \cdot t))+4 \sin t"

The final answer can be obtained after the return to the variable x:

"y=e^{-\frac{\ln(1+x)}{2}}(C_1 \cos (\sqrt{3}/2 \ln(1+x) )+ C_2 \sin(\sqrt{3}/2 \ln(1+x)))+4 \sin (\ln(1+x))"


Answer: "y=e^{-\frac{\ln(1+x)}{2}}(C_1 \cos (\sqrt{3}/2 \ln(1+x) )+ C_2 \sin(\sqrt{3}/2 \ln(1+x)))+4 \sin (\ln(1+x))"